OpenStudy (anonymous):

how to do solve: 2y -3x =7, 4Y-12=5X

OpenStudy (anonymous):

ok i got this for ur answer , y=3/2x +7/2, and y=5/2x+3 but i am not sure on the first one but the second one i know is right :)

OpenStudy (texaschic101):

2y - 3x = 7 4y - 12 = 5x --> 4y - 5x = 12 now we have : 2y - 3x = 7 -->(-2) 4y - 5x = 12 ------------- -4y + 6x = -14 (result of multiplying by -2) 4y - 5x = 12 -------------add x = -2 now sub -2 in for x in either of the equations to find y 2y - 3x = 7 2y - 3(-2) = 7 2y + 6 = 7 2y = 7 - 6 2y = 1 y = 1/2 check... 4y - 5x = 12 4(1/2) - 5(-2) = 12 2 + 10 = 12 12 = 12 (correct) x = -2 and y = 1/2